Question 2 1 pts You need to make a 500 mL buffer containing 35 mM NaCl and 50 mM Tris buffer. You already have a 5 M NaCl stock solution and a 1 M Tris stock solution. How much NaCl stock do you need (mL)?
Answer #1

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 5 M

M2 = 35 mM = 0.035 M

V2 = 500 mL

use:

M1*V1 = M2*V2

V1 = (M2 * V2) / M1

V1 = (0.035*500)/5

V1 = 3.50 mL

Answer: 3.50 mL